Please help me with this word problem: Maurice is building a 2 foot tall model of the Gateway Arch is St. Louis, Missouri. If he is using a 3 inch: 78.75 scale, how tall is the actual arch?
Answer by tommyt3rd(5050)

x/2 = 78.75/3
x=2*78.75/3=52.5